OpenStudy (anonymous):

Okai so I'm on my phone and I can't reply for some reason, but I am doing a exam on my homeschooling thing. I am almost done. I just have 2 exams left and then I'm done with school for the summer. Well I'm enfolding back in school, so I have to pass these exams because right now I'm failing Algebra. I'm not smart in math even if someone explains it to me like a million times, I'll somewhat get what you are doing if I see how you're doing it. I guess. 3x-y=-1 3x+3y=27 A. (3, 10) B. (3, 6) C. (1, 4) D. (2, 7) I'll probably post like 12 more problems after this is answered. Just saying.

OpenStudy (anonymous):

And hello again. We multiply first equation by 3(both sides) 9x - 3y = -3 3x + 3y = 27 use same sum rule again. 12x = 24 x = 24/12 = 2 3*2 + 3y = 27 6 + 3y = 27 3y = 21 y = 21/3 = 7 Check: 9 * 2 - 3 * 7 = -3 18 - 21 = -3 -3 = -3 Variant D(2 7)